# Orthogonal Howe duality and dynamical (split) symmetric pairs

### v2

@@ 4.2. ABRR equation
$[J(\lambda),1\otimes\gamma(\lambda)]=\left(\sum\limits_{\alpha\in R_{+}}F_{\alpha}\otimes E_{\alpha}\right)J(\lambda).$
@@ 5. Boundary dynamical fusion
In this section, we study a boundary analog of the dynamical fusion construction of Section 4.
@@ 5.1. Boundary fusion operator
Let $X$ be an integrable $\mathfrak{k}$-representation. According to [ReshetikhinStokman], it turns out that a boundary dynamical analog of the intertwiners of Section 4 is the space of $\mathfrak{k}$-intertwiners $\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda},X)$. For instance, the following result is similar to Theorem 4.6.
@@
Likewise, there is an analog of the fusion operator of 4.8. Namely, let $V$ be an integrable representation of $\mathfrak{sl}_{2}$. Consider a weight vector $v\in V$ and some element $\xi\in X$. The tensor product $X\otimes V$ has a natural $\mathfrak{k}$-module structure, so by Proposition 5.1 we have a canonical identification
@@
Now let $\lambda,\mu$ be generic weights. Consider a weight vector $v\in V[\lambda-\mu]$ and an element $\xi\in X$. It follows from Theorem 4.6 and Proposition 5.1 that any $\mathfrak{k}$-intertwiner in $\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda},X\otimes V)$ can be uniquely presented as the composition
@@
The next proposition is a boundary analog of Theorem 4.9.
@@
  • Using 4.3, there exists a universal invertible element $G(\lambda)$ in a certain completion of $\mathrm{U}(\mathfrak{k})\otimes\mathrm{U}(\mathfrak{b})^{\mathrm{gen}}$ whose action on any $X\otimes V$ is given by $G_{X,V}(\lambda)$.
@@
[…]
@@
To prove the second statement, let $X$ be an integrable $\mathfrak{k}$-module and $V$ a finite-dimensional $\mathfrak{g}$-representation. Recall that in the notations of 5.4, we have
@@
By Theorem 4.6, the singular vector $\Phi_{\lambda}^{v}(m_{\lambda})\in M_{\mu}\otimes V$ can be computed using the universal element $\Xi(\lambda)$ in a completion of $\mathrm{U}(\mathfrak{n}_{-})\otimes\mathrm{U}(\mathfrak{b}_{+})^{\mathrm{gen}}$. Observe that $\phi^{\xi}_{\mu}$ comes from the action of the universal homomorphism given by the composition
@@
Let $\mathfrak{g}=\mathfrak{sl}_{2}$. According to Example 4.7, the boundary fusion operator is given by
@@
where we use the normal ordering on the first tensor component: the operator $h|_{V}$ acts first. Using Example 5.3, we obtain
@@
Another example is when $X=\mathbb{C}$ is the trivial representation of $\mathfrak{k}$. Example 5.3 implies that
@@
where $(2k)!!$! is the double factorial. Observe its similarity to the quantum $K$-matrix in rank one [DobsonKolb, (3.39)].
@@ 5.2. Boundary ABRR equation
The boundary fusion operator can be constructed inductively using an analog of the ABRR equation of Theorem 4.13; in what follows, we will use the notations of the theorem.
@@
(5.6) $[G(\lambda),1\otimes\gamma(\lambda)]=\left(\sum\limits_{\alpha\in R_{+}}B_{\alpha}\otimes E_{\alpha}-1\otimes E_{\alpha}^{2}\right)G(\lambda).$
@@
After multiplying this equality by $-\frac{1}{2}$, rearrangement and simplification of terms yields
@@
Applying both sides of equation 5.10 to $m_{\lambda}$, we conclude that
$[G_{X,V}(\lambda),1\otimes\gamma(\lambda)]\cdot\xi\otimes v=\sum_{\alpha\in R_{+}}(B_{\alpha}\otimes E_{\alpha}-1\otimes E_{\alpha}^{2})\cdot G_{X,V}(\lambda)\cdot\xi\otimes v,$
@@
for some polynomials $q_{k}(b,\lambda,h^{(2)})$ in $b\in\mathrm{U}(\mathfrak{k})$ with values in $\mathcal{M}(\mathfrak{h}^{*})$ such that $q_{0}=1$, and $h^{(2)}$ is the weight operator on the second tensor component. Proceeding as in Example 4.14, we derive the following equation:
@@ 6. Background on dynamical Weyl group
The dynamical fusion operators give rise to a dynamical quantum group (Hopf algebroid) [EtingofVarchenkoExchange]. It is observed in [EtingofVarchenkoDynamicalWeyl] that there is a dynamical Weyl group which is analogous to the quantum Weyl group for the usual Drinfeld–Jimbo quantum group. In this section, we recall the construction of the dynamical Weyl group operators. In what follows, we use definitions and notations of Section 4. In particular, $\mathfrak{g}$ is a simple Lie algebra.
@@ 6.1. Dynamical Weyl group
As in Section 4, let $V$ be a finite-dimensional $\mathfrak{g}$-representation, and consider the space $\mathrm{Hom}_{\mathfrak{g}}(M_{\lambda},M_{\mu}\otimes V)$. Observe that the intertwiner $\Phi_{\lambda}^{v}$ from Theorem 4.6 is defined for all sufficiently dominant integral weights.
@@
[CherednikQKZ, Definition 2.1] Consider a collection of operators $\{G_{\alpha}(\mu)\}$ for $\alpha\in R_{+}$ and $\mu\in\mathbb{C}$. Define the functions $G^{\alpha}:=G^{\alpha}(\langle\lambda,\alpha^{\vee}\rangle)$ on $\mathfrak{h}^{*}$. Such a collection is called a closed $R$-matrix if, for every pair of roots $\alpha,\beta$, they satisfy one of the following relations:
@@ 7.1. Boundary dynamical Weyl group: $\mathfrak{so}_{2}\subset\mathfrak{sl}_{2}$
In this subsection, we construct a boundary analog of the dynamical Weyl group as in Section 6.
@@
Let $X$ be an integrable $\mathfrak{so}_{2}$-representation. Consider $\xi\in X$ and the corresponding $\mathfrak{so}_{2}$-intertwiner $\phi_{\lambda}^{\xi}\in\mathrm{Hom}_{\mathfrak{so}_{2}}(M_{\lambda},X)$ as in Proposition 5.1. Assume that $\langle\lambda,\alpha^{\vee}\rangle\in\mathbb{Z}_{\geq 0}$. Define $A_{X}^{\mathfrak{so}_{2}}(\lambda)\in\mathrm{End}_{\mathbb{C}}(X)$ by
@@
According to Example 5.3, the action of $A_{\mathbb{C}_{x}}(\lambda)$ is given by the polynomial $p_{\lambda+1}(x,\lambda)$. It is clear that its degree is $\lambda+1$ and that the coefficient in front of $x^{\lambda+1}$ is equal to $1/(\lambda+1)!$. Therefore, it is enough to prove that the roots of $p_{\lambda+1}(x,\lambda)$ are
@@
[…]
@@
Since $b=f-e$ has eigenvalues $i$ and $-i$ in the natural representation $V_{1}\cong\mathbb{C}^{2}$, $b$ is conjugate to $ih$ inside $\mathfrak{sl}_{2}$ and therefore acts semisimply in the finite dimensional module $V_{\lambda}$ with the eigenvalues (7.3). Thus, for $x$ in the set (7.3), we have $\dim\mathrm{Hom}_{\mathfrak{k}}(V_{\lambda},\mathbb{C}_{x})=1$. Also, from Proposition 5.1 we have $\dim\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda},\mathbb{C}_{x})=\dim\mathbb{C}_{x}=1$. By exactness, $\pi\circ(-)$ is injective, and therefore is surjective. ∎
@@
[…]
@@ 7.2. Divided power formula
The following theorem is a dynamical analog of Proposition 3.10.
@@
It follows from equation (7.6) and 7.8 that
@@ 7.3. Boundary dynamical Weyl group: $\mathfrak{k}\subset\mathfrak{g}$
as in 7.7 for all roots, not necessarily simple.
@@
for some rational functions $f_{k}(\lambda)$. Moreover, by integrability and (7.6), we can choose the collection $\{x_{k}\}$ such that $x_{k}$’s are pairwise non-congruent modulo $2\sqrt{-1}$ for every $k$. Since these equalities are satisfied for all sufficiently positive integral $\lambda$, they are satisfied for all $\lambda$ by Proposition A.2. ∎
@@
Let $\mathfrak{g}=\mathfrak{sl}_{2}$, with standard generators $\{e,h,f\}$, and $\mathfrak{k}=\mathfrak{so}_{2}=\mathrm{span}(b)$, where $b:=f-e$. Denote by $\{v_{+},v_{-}\}$ the standard $h$-weight basis of $V(1)$.
The eigenvectors of $b$ are $v_{+}-iv_{-}$, with eigenvalue $i$, and $v_{+}+iv_{-}$, with eigenvalue $-i$. Therefore, the matrix of $A^{\mathfrak{so}_{2}}_{\mathbb{C}_{x}\otimes V(1)}(\lambda)$ in the standard basis is
@@
We use Example 5.6 and Example 6.1 to compute that the matrix of
@@
Thanks to 7.8, the two matrices coincide.
The following theorem is the boundary analog of Proposition 6.5.
@@
Rewriting Proposition 7.16 gives that
@@
Therefore, it is enough to prove the braid relations for irreducible representations $X_{1},\ldots,X_{n}$ such that any integrable representation of $\mathfrak{k}$ is a direct summand of $X_{i}\otimes V$ for some $i$ and a suitable representation $V$ of $\mathfrak{g}$. Note that it is enough to prove this when $\mathfrak{g}$ is a rank 2 algebra. This is done in Appendix B. ∎
@@ 7.4. Boundary $B$-operators
As in Section 6, we can define the $B$-operators. Recall from 3.6 that the boundary Weyl group operator is defined by $\tilde{s}^{\mathfrak{k}}_{\alpha}=\exp\left({\frac{\pi b_{\alpha}}{2}}\right)$. Since $\sin(x)=(\exp(ix)-\exp(-ix))/{2i}$, we find that
@@
The shift of variable by $-1$ corresponds to the $\rho$-shift. It differs from the $B$-operators (6.3) for $\mathfrak{g}$, but it will be more convenient for the $(\mathfrak{so}_{2n},\mathrm{O}_{m})$-duality of Section 10.5.
@@
Thanks to the next lemma, these operators behave well with respect to the boundary Weyl group operators of Section 3.
@@
We want to compute the action of conjugation by $\tilde{s}_{i}^{\mathfrak{k}}$, the adjoint action of $\tilde{s}_{i}^{\mathfrak{k}}$, on the $B$ operator associated to $\beta$. By 3.15, we obtain that $\mathrm{Ad}_{\tilde{s}^{\mathfrak{k}}_{i}}(b_{\beta})=\pm b_{s_{i}(\beta)}$. Application of $\mathrm{Ad}_{\tilde{s}^{\mathfrak{k}}_{i}}$ to the boundary $B$-operator associated to $\beta$ then gives
@@
Follows from equation (7.18) and 7.22. ∎
@@
It is enough to prove the lemma for rank 2 algebras which is done case-by-case in Appendix B. ∎
The following theorem is a boundary analog of Theorem 6.9.
@@
The operators $\{\mathcal{B}^{\mathfrak{k}}_{\beta}(\lambda)\}$ form a closed $R$-matrix in the sense of 6.8.
@@
Observe the relations in 6.8 are equivalent to the same relations for $G_{\alpha}^{-1}$. So, it is enough to prove the statement for
@@
It follows easily from Theorem 7.18 that the operators $\tilde{A}^{\mathfrak{k},\alpha}(\lambda)^{-1}:=A^{\mathfrak{k},\alpha}(\lambda-\rho)^{-1}$ satisfy the unshifted dynamical braid relation
@@
and the same is true for the sine operators $S_{\alpha}(\lambda)$ by 7.24. Multiply both sides by
@@
on the right. By pushing the sine terms to the left accordingly and using 7.23, we conclude. ∎
@@
Note that the relation between $B$-operators and dynamical Weyl group is different in the boundary case compared to the one for $\mathfrak{g}$ of Theorem 6.7. Theorem 7.25 was the main motivation to use the sine factor instead of the boundary Weyl group operator.

### v3

@@ 4.2. ABRR equation
$J(\lambda),1\otimes\gamma(\lambda)]=\left(\sum\limits_{\alpha\in R_{+}}F_{\alpha}\otimes E_{\alpha}\right)J(\lambda).$
@@ 5. Boundary dynamical fusion
In this section, we study a boundary analog of the dynamical fusion construction of Section˜4.
@@ 5.1. Boundary fusion operator
Let $X$ be an integrable $\mathfrak{k}$-representation. According to [ReshetikhinStokman], it turns out that a boundary dynamical analog of the intertwiners of Section˜4 is the space of $\mathfrak{k}$-intertwiners $\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda},X)$. For instance, the following result is similar to Theorem˜4.6.
@@
Likewise, there is an analog of the fusion operator of ˜4.8. Namely, let $V$ be an integrable representation of $\mathfrak{sl}_{2}$. Consider a weight vector $v\in V$ and some element $\xi\in X$. The tensor product $X\otimes V$ has a natural $\mathfrak{k}$-module structure, so by Proposition˜5.1 we have a canonical identification
@@
Now let $\lambda,\mu$ be generic weights. Consider a weight vector $v\in V[\lambda-\mu]$ and an element $\xi\in X$. It follows from Theorem˜4.6 and Proposition˜5.1 that any $\mathfrak{k}$-intertwiner in $\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda},X\otimes V)$ can be uniquely presented as the composition
@@
The next proposition is a boundary analog of Theorem˜4.9.
@@
  • Using ˜4.3, there exists a universal invertible element $G(\lambda)$ in a certain completion of $\mathrm{U}(\mathfrak{k})\otimes\mathrm{U}(\mathfrak{b})^{\mathrm{gen}}$ whose action on any $X\otimes V$ is given by $G_{X,V}(\lambda)$.
@@
${M_{\lambda}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\lambda}^{w}}$${M_{\mu}\otimes W\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\mu}^{G_{X,V}(\mu)(\xi\otimes v)}\otimes\mathrm{id}_{W}}$$\scriptstyle{\Phi_{\mu}^{v}\otimes\mathrm{id}_{W}}$${M_{\nu}\otimes V\otimes W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\nu}^{\xi}\otimes\mathrm{id}_{W}}$${X\otimes V\otimes W}$
@@
To prove the second statement, let $X$ be an integrable $\mathfrak{k}$-module and $V$ a finite-dimensional $\mathfrak{g}$-representation. Recall that in the notations of ˜5.4, we have
@@
By Theorem˜4.6, the singular vector $\Phi_{\lambda}^{v}(m_{\lambda})\in M_{\mu}\otimes V$ can be computed using the universal element $\Xi(\lambda)$ in a completion of $\mathrm{U}(\mathfrak{n}_{-})\otimes\mathrm{U}(\mathfrak{b}_{+})^{\mathrm{gen}}$. Observe that $\phi^{\xi}_{\mu}$ comes from the action of the universal homomorphism given by the composition
@@
Let $\mathfrak{g}=\mathfrak{sl}_{2}$. According to Example˜4.7, the boundary fusion operator is given by
@@
where we use the normal ordering on the first tensor component: the operator $h|_{V}$ acts first. Using Example˜5.3, we obtain
@@
Another example is when $X=\mathbb{C}$ is the trivial representation of $\mathfrak{k}$. Example˜5.3 implies that
@@
where $(2k)!!$ is the double factorial. Observe its similarity to the quantum $K$-matrix in rank one [DobsonKolb, (3.39)].
@@ 5.2. Boundary ABRR equation
The boundary fusion operator can be constructed inductively using an analog of the ABRR equation of Theorem˜4.13; in what follows, we will use the notations of the theorem.
@@
(5.6) $G(\lambda),1\otimes\gamma(\lambda)]=\left(\sum\limits_{\alpha\in R_{+}}B_{\alpha}\otimes E_{\alpha}-1\otimes E_{\alpha}^{2}\right)G(\lambda).$
@@
After multiplying this equality by $\frac{1}{2}$, rearrangement and simplification of terms yields
@@
Applying both sides of equation (5.10) to $m_{\lambda}$, we conclude that
$G_{X,V}(\lambda),1\otimes\gamma(\lambda)]\cdot\xi\otimes v=\sum_{\alpha\in R_{+}}(B_{\alpha}\otimes E_{\alpha}-1\otimes E_{\alpha}^{2})\cdot G_{X,V}(\lambda)\cdot\xi\otimes v,$
@@
for some polynomials $q_{k}(b,\lambda,h^{(2)})$ in $b\in\mathrm{U}(\mathfrak{k})$ with values in $\mathcal{M}(\mathfrak{h}^{*})$ such that $q_{0}=1$, and $h^{(2)}$ is the weight operator on the second tensor component. Proceeding as in Example˜4.14, we derive the following equation:
@@ 6. Background on dynamical Weyl group
The dynamical fusion operators give rise to a dynamical quantum group (Hopf algebroid) [EtingofVarchenkoExchange]. It is observed in [EtingofVarchenkoDynamicalWeyl] that there is a dynamical Weyl group which is analogous to the quantum Weyl group for the usual Drinfeld–Jimbo quantum group. In this section, we recall the construction of the dynamical Weyl group operators. In what follows, we use definitions and notations of Section˜4. In particular, $\mathfrak{g}$ is a simple Lie algebra.
@@ 6.1. Dynamical Weyl group
As in Section˜4, let $V$ be a finite-dimensional $\mathfrak{g}$-representation, and consider the space $\mathrm{Hom}_{\mathfrak{g}}(M_{\lambda},M_{\mu}\otimes V)$. Observe that the intertwiner $\Phi_{\lambda}^{v}$ from Theorem 4.6 is defined for all sufficiently dominant integral weights.
@@
[CherednikQKZ, Definition 2.1] Let $V$ be a vector space. Consider a collection $\{G_{\alpha}(\mu)\}_{\alpha\in R^{+},\mu\in\mathbb{C}}$ of operators on $V$. Define functions $G^{\alpha}:\mathfrak{h}^{*}\rightarrow\mathrm{End}(V)$ by $G^{\alpha}(\lambda):=G_{\alpha}(\langle\lambda,\alpha^{\vee}\rangle)$. Such a collection is called a closed $R$-matrix if, for every pair of roots $\alpha,\beta$, they satisfy one of the following relations:
@@ 7.1. Boundary dynamical Weyl group: $\mathfrak{so}_{2}\subset\mathfrak{sl}_{2}$
In this subsection, we construct a boundary analog of the dynamical Weyl group as in Section˜6.
@@
Let $X$ be an integrable $\mathfrak{so}_{2}$-representation. Consider $\xi\in X$ and the corresponding $\mathfrak{so}_{2}$-intertwiner $\phi_{\lambda}^{\xi}\in\mathrm{Hom}_{\mathfrak{so}_{2}}(M_{\lambda},X)$ as in Proposition˜5.1. Assume that $\langle\lambda,\alpha^{\vee}\rangle\in\mathbb{Z}_{\geq 0}$. Define $A_{X}^{\mathfrak{so}_{2}}(\lambda)\in\mathrm{End}_{\mathbb{C}}(X)$ by
@@
According to Example˜5.3, the action of $A_{\mathbb{C}_{x}}(\lambda)$ is given by the polynomial $p_{\lambda+1}(x,\lambda)$. It is clear that its degree is $\lambda+1$ and that the coefficient in front of $x^{\lambda+1}$ is equal to $1/(\lambda+1)!$. Therefore, it is enough to prove that the roots of $p_{\lambda+1}(x,\lambda)$ are
@@
$0\rightarrow M_{s\cdot\lambda}\xrightarrow{\iota_{s\cdot\lambda}^{\lambda}}M_{\lambda}\xrightarrow{\pi}V_{\lambda}\rightarrow 0,$
@@
Since b=feb=f-e has eigenvalues $i$ and i-i in the natural representation $V_{1}\cong\mathbb{C}^{2}$, $b$ is conjugate to $ih$ inside $\mathfrak{sl}_{2}$ and therefore acts semisimply in the finite dimensional module $V_{\lambda}$ with the eigenvalues (7.3). Thus, for $x$ in the set (7.3), we have $\dim\mathrm{Hom}_{\mathfrak{k}}(V_{\lambda},\mathbb{C}_{x})=1$. Also, from Proposition 5.1 we have $\dim\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda},\mathbb{C}_{x})=\dim\mathbb{C}_{x}=1$. By exactness, $\pi\circ(-)$ is injective, and therefore is surjective. ∎
@@
Euler’s gamma function $\Gamma(z)$ is defined as the analytic continuation of the function $z\mapsto\int_{0}^{\infty}t^{z-1}\exp(-t)dt$, for $z\in\mathbb{R}_{\geq 0}$, to a meromorphic function on $\mathbb{C}$.
@@ 7.2. Divided power formula
The following theorem is a dynamical analog of Proposition˜3.10.
@@
It follows from equation (7.6) and ˜7.8 that
@@ 7.3. Boundary dynamical Weyl group: $\mathfrak{k}\subset\mathfrak{g}$
as in ˜7.7 for all roots, not necessarily simple.
@@
for some rational functions $f_{k}(\lambda)$. Moreover, by integrability and (7.6), we can choose the collection $\{x_{k}\}$ such that $x_{k}$’s are pairwise non-congruent modulo $2\sqrt{-1}$ for every $k$. Since these equalities are satisfied for all sufficiently positive integral $\lambda$, they are satisfied for all $\lambda$ by Proposition˜A.2. ∎
@@
Let $\mathfrak{g}=\mathfrak{sl}_{2}$, with standard generators $\{e,h,f\}$, and $\mathfrak{k}=\mathfrak{so}_{2}=\mathrm{span}(b)$, where b:=feb:=f-e. Denote by $\{v_{+},v_{-}\}$ the standard $h$-weight basis of $V(1)$.
The eigenvectors of $b$ are $v_{+}-iv_{-}$, with eigenvalue $i$, and $v_{+}+iv_{-}$, with eigenvalue i-i. Therefore, the matrix of $A^{\mathfrak{so}_{2}}_{\mathbb{C}_{x}\otimes V(1)}(\lambda)$ in the standard basis is
@@
We use Example˜5.6 and Example˜6.1 to compute that the matrix of
@@
Thanks to ˜7.8, the two matrices coincide.
The following theorem is the boundary analog of Proposition˜6.5.
@@
Rewriting Proposition˜7.16 gives that
@@
Therefore, it is enough to prove the braid relations for irreducible representations $X_{1},\ldots,X_{n}$ such that any integrable representation of $\mathfrak{k}$ is a direct summand of $X_{i}\otimes V$ for some $i$ and a suitable representation $V$ of $\mathfrak{g}$. Note that it is enough to prove this when $\mathfrak{g}$ is a rank 2 algebra. This is done in Appendix˜B. ∎
@@ 7.4. Boundary $B$-operators
As in Section˜6, we can define the $B$-operators. Recall from ˜3.6 that the boundary Weyl group operator is defined by $\tilde{s}^{\mathfrak{k}}_{\alpha}=\exp\left({\frac{\pi b_{\alpha}}{2}}\right)$. Since $\sin(x)=(\exp(ix)-\exp(-ix))/{2i}$, we find that
@@
The shift of variable by 1-1 corresponds to the $\rho$-shift. It differs from the $B$-operators (6.3) for $\mathfrak{g}$, but it will be more convenient for the $(\mathfrak{so}_{2n},\mathrm{O}_{m})$-duality of Section˜10.5.
@@
Thanks to the next lemma, these operators behave well with respect to the boundary Weyl group operators of Section˜3.
@@
We want to compute the action of conjugation by $\tilde{s}_{i}^{\mathfrak{k}}$, the adjoint action of $\tilde{s}_{i}^{\mathfrak{k}}$, on the $B$ operator associated to $\beta$. By ˜3.15, we obtain that $\mathrm{Ad}_{\tilde{s}^{\mathfrak{k}}_{i}}(b_{\beta})=\pm b_{s_{i}(\beta)}$. Application of $\mathrm{Ad}_{\tilde{s}^{\mathfrak{k}}_{i}}$ to the boundary $B$-operator associated to $\beta$ then gives
@@
Follows from equation (7.18) and ˜7.22. ∎
@@
It is enough to prove the lemma for rank 2 algebras which is done case-by-case in Appendix˜B. ∎
The following theorem is a boundary analog of Theorem˜6.9.
@@
The operators $\{\mathcal{B}^{\mathfrak{k}}_{\beta}(\lambda)\}$ form a closed $R$-matrix in the sense of ˜6.8.
@@
Observe the relations in ˜6.8 are equivalent to the same relations for $G_{\alpha}^{-1}$. So, it is enough to prove the statement for
@@
It follows easily from Theorem˜7.18 that the operators $\tilde{A}^{\mathfrak{k},\alpha}(\lambda)^{-1}:=A^{\mathfrak{k},\alpha}(\lambda-\rho)^{-1}$ satisfy the unshifted dynamical braid relation
@@
and the same is true for the sine operators $S_{\alpha}(\lambda)$ by ˜7.24. Multiply both sides by
@@
on the right. By pushing the sine terms to the left accordingly and using ˜7.23, we conclude. ∎
@@
Note that the relation between $B$-operators and dynamical Weyl group is different in the boundary case compared to the one for $\mathfrak{g}$ of Theorem˜6.7. Theorem˜7.25 was the main motivation to use the sine factor instead of the boundary Weyl group operator.